OpenStudy (anonymous):

use the binomial theorem to expand the binomial. (3v+s)^5

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

Possible answers : s5 + 45s4v + 270s3v2 + 810s2v3 + 1215sv4 + 729v5 s5 + 15s4v + 90s3v2 + 270s2v3 + 405sv4 + 243v5 s5 + 15s4v + 90s3 + 270s2 + 405s + 243 s5 – 5s4v + 10s3v2 – 10s2v3 + 5sv4 – v5

OpenStudy (anonymous):

@Butterfly16

OpenStudy (campbell_st):

ok... so pascal's triangle will give the coefficients  starting with (3v)^5 and (s)^0 decrease the power or 3v by 1 each time and increase the power of s so you will have, using the last line of pascals triangle.. \[1 \times (3v)^5\times (s)^0 + 5\times (3v)^4\times(s)^1 + 10 \times (3v)^3 \times (s)^2 +...... \] you need to continue until your have (3v)^0 and (s)^5